# From Word to World: Evaluate and Mitigate Culture Bias via Word Association Test

### v1

@@ Abstract
The human-centered word association test (WAT) serves as a cognitive proxy, revealing sociocultural variations through lexical-semantic patterns. We extend this test into an LLM-adaptive, free-relation task to assess the alignment of large language models (LLMs) with cross-cultural cognition. To mitigate the culture preference, we propose CultureSteer, an innovative approach that integrates a culture-aware steering mechanism to guide semantic representations toward culturally specific spaces. Experiments show that current LLMs exhibit significant bias toward Western cultural (notably in American) schemas at the word association level. In contrast, our model substantially improves cross-cultural alignment, surpassing prompt-based methods in capturing diverse semantic associations. Further validation on culture-sensitive downstream tasks confirms its efficacy in fostering cognitive alignment across cultures. This work contributes a novel methodological paradigm for enhancing cultural awareness in LLMs, advancing the development of more inclusive language technologies.
From Word to World:
Evaluate and Mitigate Culture Bias via Word Association Test
Xunlian Dai1,2, Li Zhou1thanks: Corresponding author, Benyou Wang1, Haizhou Li1,3 1The Chinese University of Hong Kong, Shenzhen 2Wuhan University of Technology 3Shenzhen Research Institute of Big Data, lizhou21@cuhk.edu.cn
@@ 1 Introduction
Word associations, often perceived as commonplace neural activities Pulvermüller (1999); Anderson et al. (2017); Khanna et al. (2024), are deeply rooted in lived environments and personal experiences. The neurocognitive research Schneider et al. (2024) find out that transient human associations reflect not only situational factors but also ingrained in sociocultural identities. Cross-cultural WAT also unveil how culture modulates perceptual and interactive patterns Szalay and Deese (2024).
@@
Multilingual LLMs exhibit persistent cultural biases, ranging from stereotypical associations Abid et al. (2021); Bano et al. (2025) to value misalignments Masoud et al. (2023); Cao et al. (2023a); Jiang et al. (2024). Prior research has largely concentrated on addressing these biases by aligning multicultural cognition through prompt-based methods Choenni and Shutova (2024); Wang et al. (2024); Sato et al. (2024), which depend on explicitly providing cultural context to elicit targeted responses. On the other hand, fine-tuning approaches Xu et al. (2024); Yao et al. (2025) are knowledge-driven but still require explicit cultural settings during inference. In contrast, our approach implicitly enhances the model’s cultural awareness and sensitivity by simulating human word association tasks and learning semantic spaces of cultural preferences within the model itself, directly capturing latent cultural relationships. Our contributions can be summarized as follows:
@@ 2 Related Work
Extensive research has investigated cultural biases in LLMs across domains such as value preference Benkler et al. (2023); Cao et al. (2023a); Jiang et al. (2024); Durmus et al. (2023), knowledge perception Palta and Rudinger (2023); Shen et al. (2024), and moral measurement Jinnai (2024); Rao et al. (2023, 2024), revealing persistent Western-centric value preferences and regionally divergent commonsense knowledge. To improve cultural perception, current approaches predominantly employ prompt engineering Wang et al. (2024); AlKhamissi et al. (2024); Zhou et al. (2025); Masoud et al. (2025), explicitly embedding cultural context in prompts to evoke culturally sensitive responses. Alternatively, fine-tuning methods Yao et al. (2025); Xu et al. (2024) have been proposed to enhance cultural alignment via few-shot learning or cross-cultural similarity modeling. Our approach stands out by proposing a novel word-level perspective through free word associations, directly integrating cultural knowledge into the model’s internal representations via associative fine-tuning, thereby achieving deeper cultural alignment beyond prompt-based methods.
@@
Recent work adopts word association tasks to evaluate the capabilities of LLMs across various domains. For instance, uncovering gender stereotypes Abramski et al. (2024b) and assessing color associative perception Fukushima et al. (2024). These studies primarily focus on word association mappings under specific relationships, emphasizing explicit word generation and partially relying on manual interpretation during evaluation Abramski et al. (2024a); Vintar et al. (2024).
@@ 3.1 Human-centered task
The WAT for human participants is conceptualized as an open-ended activity in a free-relation format, allowing for flexible and diverse associations that more accurately reflect human cognitive processes. [^1] https://smallworldofwords.org Before conducting the WAT, demographic information, such as age, gender, education, native language, and region, is collected for each participant. Given a cue word $w$, the participant $u$ is asked to provide the first word that comes to mind and may optionally add a second and third word sequentially, resulting in up to three associated words, represented as $u_{w}^{c}=\left\{a_{1}^{u},a_{2}^{u},a_{3}^{u}\right\}$, where $c$ denotes the participant’s cultural background. Within this cultural context, the collective set of associative words for a given cue word $w$, aggregated across all participants $U_{w}^{c}=\left\{u_{w}^{c}\right\}$, is represented as $A_{w}^{c}=\left\{\left(a_{1},f_{1}\right),\cdots,\left(a_{n},f_{n}\right)\right\}$. Here, $f_{i}\geqslant f_{i+1}$, and $f_{i}$ represents the frequency with which the associative word $a_{i}$ appears in $U_{w}^{c}$. Across different cultural contexts, the associative words in $A_{w}^{c}$ may exhibit nuanced variations in both semantic content and frequency distribution.
@@ 3.2 LLM-adaptive task
Directly employing the human-centered task paradigm to conduct WAT with LLMs proves insufficient for effectively capturing variations across entire cultural groups. Inspired by Fierro et al. (2024) and Zhou et al. (2025), we adapt this task into a word prediction framework, designed as a retrieval process for candidate associative words. Specifically, we employ a template $q$, such as “When {cue_word} is mentioned, people often associate it with the following words:”, to prompt the LLM and elicit its predictions of associated words. The input is represented as $x=q(w)$, and the output is denoted as $o=M(q(w))$, where the generalized function $M(\cdot)$ represents the inference process of the LLM.
@@ 3.3 Evaluation metric
$\text{R@K}&={\frac{\sum_{i=1}^{K}{\mathbb{I}\left(\hat{a}_{i}\in A_{w}^{c}\right)}}{N}}\\ &=\frac{\sum_{i=1}^{N}{\mathbb{I}\left(a_{i}\in\text{Top-K}\left(\hat{A}_{w}^{c}\right)\right)}}{N}\end{aligned}$ (2)
@@ 4 CultureSteer: Modeling and Controlling Cultural Awareness in LLMs
Inspired by the findings of LM-Steer Han et al. (2024), which suggest that, with certain assumptions, shifting styles in language models is equivalent to a linear transformation in the word embedding space, we extend this perspective to the cultural domain. Specifically, we hypothesize that different cultures define distinct semantic association spaces shaped by cultural preferences, which can be modeled through culture-specific linear transformations within the embedding space.
Based on this, we introduce the CultureSteer model, which leverages both human-centered and LLM-adaptive WAT tasks to (1) equip LLMs with human-like word association perception, thereby enhancing their cognitive abilities, and (2) enable the word association capabilities of LLMs to align across different cultures. The overall architecture of the model is illustrated in Figure 2.
@@
PWR@K K=3 K=5 K=10 K=20 PWR@K K=3 K=5 K=10 K=20
Baseline 12.37 19.57 29.25 42.54 Baseline 15.22 19.05 26.18 35.90
CSP 30.30 37.61 48.95 60.95 CSP 20.34 28.19 39.05 50.65
CCT 25.92 32.98 44.05 55.54 CCT 22.59 30.51 42.31 54.76
USA CultureSteer 38.14 50.30 67.58 80.13 CultureSteer 36.34 46.06 59.78 72.05
Baseline 10.69 16.53 24.96 35.92 Baseline 12.43 16.16 22.05 29.97
CSP 24.31 30.64 39.91 49.27 CSP 16.43 22.14 30.58 39.86
CCT 21.58 27.74 35.70 44.96 CCT 18.17 24.27 33.54 42.72
UK CultureSteer 29.19 38.84 51.45 61.88 CultureSteer 28.23 34.88 46.10 57.53
Baseline 9.53 15.62 23.25 33.67 Baseline 11.87 15.07 21.35 28.41
CSP 23.17 28.62 36.78 45.88 CSP 15.58 21.21 29.02 37.60
CCT 20.21 25.64 33.37 42.41 CCT 17.16 22.67 31.61 40.63
OC CultureSteer 27.56 35.44 47.65 58.77 CultureSteer 25.05 32.22 41.55 52.22
Baseline 7.31 11.03 18.43 24.01 Baseline 8.61 11.65 17.16 24.11
CSP 10.52 15.83 23.34 29.81 CSP 15.56 21.59 31.77 42.81
CCT 9.71 14.08 21.02 28.00 CCT 12.32 17.71 26.57 36.85
CN CultureSteer 12.77 19.04 27.26 33.53 CultureSteer 21.78 30.13 44.88 58.76
Baseline 9.98 15.69 23.97 34.04 Baseline 12.03 15.48 21.69 29.60
CSP 22.08 28.18 37.28 46.48 CSP 16.98 23.28 32.61 42.73
CCT 19.36 25.11 33.54 42.73 CCT 17.56 23.79 33.51 43.74
Average CultureSteer 26.92 35.91 48.49 58.58 CultureSteer 27.85 35.82 48.08 60.14
@@ 5.1 Dataset Creation
We utilized the officially released datasets from the Small World of Words Project (SWOW) [^4] Designed as a human-centered Word Association Task (WAT).. For our study, we selected data in two languages: English and Mandarin Chinese, referred to as SWOW-EN De Deyne et al. (2019) and SWOW-ZH Li et al. (2024), respectively. The SWOW-EN dataset comprises over 3 million responses collected from more than 90,000 participants across various English-speaking countries or regions, covering more than 12,000 cues. Moreover, the SWOW-ZH dataset contains over 2 million responses corresponding to 10,192 cues, with participants primarily from the Chinese mainland.
@@
To ensure sufficient participant representation across different cultures, we first filter the three most prominent cultural groups from SWOW-EN: the United States (USA), the United Kingdom (UK), and Oceania (OC) [^5] OC includes participant responses from Australia and New Zealand, which are combined due to their relatively small participant numbers, geographical proximity, and shared historical background.. Additionally, SWOW-ZH serves as the representative for China (CN). As a result, we construct datasets representing four distinct cultural groups: USA, UK, OC, and CN.
@@
To ensure fairness in evaluating word association perception across different cultures, we utilize the Intercontinental Dictionary Series (IDS) jung Yu and Wang (2023) to align the cue words between English and Chinese, ultimately retaining 881 cue words that are common across all four cultural groups. Given the substantial differences in participant numbers across cultures, which lead to significant variation in the number of associated words ($A_{w}^{c}$), we standardize the $A_{w}^{c}$ lists by truncating them to the size of the culture with the fewest associated words. Using the IDS framework, we categorize these cue words into 22 semantic categories based on their conceptual meanings. Examples of the final preprocessed data and statistical details are provided in Appendix B.2, Table 4 and Table 5. Finally, we split the cue words in each semantic category into 70% for training and 30% for testing.
@@ 5.2 Experimental Setup
Figure 3: Fine-grained performance comparison across 22 semantic classes in the test set with PWR@20. Red denotes Global Knowledge, green denotes Perceptual Experience, and blue denotes Cultural Ideologies. Other PWR@K (K=3, 5, 10) results are shown in Appendix D.
We conduct experiments using Llama3.1-8B Grattafiori et al. (2024) and Qwen2.5-7B Team (2024) to simulate the WAT. First, we evaluate the base versions of these models on word association tasks using a base prompt without any cultural conditioning, aiming to assess their human-like associative cognition and the default cultural representation they exhibit. Subsequently, we build CultureSteer based on these two models to enhance their cross-cultural perception capabilities. For the LLM backbone, we adopt LoRA Hu et al. (2022), a parameter-efficient tuning method, implemented via LLaMAFactory, and train the models on a single A100 GPU. Furthermore, we introduce two explicit culture-aware prompting strategies as compared methods: Culture-Specific Prompt (CSP) Zhou et al. (2025), which defines a specific cultural context in the prompt, and Cross-Culture Thinking (CCT) Xu et al. (2025), which prompts LLMs to consider the relationships between the given culture and other cultures. Detailed prompt settings are provided in Appendix B.1.

### v2

# From Word to World: Evaluate and Mitigate Culture Bias in LLMs via Word Association Test

@@ Abstract
The human-centered word association test (WAT) serves as a cognitive proxy, revealing sociocultural variations through culturally shared semantic expectations and implicit linguistic patterns shaped by lived experiences. We extend this test into an LLM-adaptive, free-relation task to assess the alignment of large language models (LLMs) with cross-cultural cognition. To address culture preference, we propose CultureSteer, an innovative approach that moves beyond superficial cultural prompting by embedding cultural-specific semantic associations directly within the model’s internal representation space. Experiments show that current LLMs exhibit significant bias toward Western (notably American) schemas at the word association level. In contrast, our model substantially improves cross-cultural alignment, capturing diverse semantic associations. Further validation on culture-sensitive downstream tasks confirms its efficacy in fostering cognitive alignment across cultures. This work contributes a novel methodological paradigm for enhancing cultural awareness in LLMs, advancing the development of more inclusive language technologies. [^1] Our code and resources are publicly available at https://github.com/hlt-cuhksz/CultureSteer.
From Word to World: Evaluate and Mitigate Culture Bias
in LLMs via Word Association Test
Xunlian Dai1, Li Zhou1thanks: Corresponding author, Benyou Wang1, Haizhou Li1,2 1The Chinese University of Hong Kong, Shenzhen 2Shenzhen Research Institute of Big Data 225040015@link.cuhk.edu.cn; lizhou21@cuhk.edu.cn
@@ 1 Introduction
Word associations, often perceived as commonplace neural activities Pulvermüller (1999); Anderson et al. (2017); Khanna et al. (2024), are deeply rooted in lived environments and personal experiences. The neurocognitive research Schneider et al. (2024) find out that transient human associations reflect not only situational factors but also ingrained in sociocultural identities. Multilingual and cross-cultural WAT also unveil how culture modulates perceptual and interactive patterns Szalay and Deese (2024); Garimella et al. (2017, 2016).
@@
Multilingual LLMs exhibit persistent cultural biases, ranging from stereotypical associations Abid et al. (2021); Bano et al. (2025) to value misalignments Masoud et al. (2023); Cao et al. (2023a); Jiang et al. (2024). Prior research has largely concentrated on addressing these biases by aligning multicultural cognition through prompt-based methods Choenni and Shutova (2024); Wang et al. (2024); Sato et al. (2024), which depend on explicitly providing cultural context to elicit targeted responses. On the other hand, fine-tuning approaches Xu et al. (2024); Yao et al. (2025) are knowledge-driven but still require explicit cultural settings during inference.
Building on this perspective, our approach follows the opinion that cultural competence in NLP is often best evaluated through diagnostic tasks that reveal latent cultural patterns Zhou et al. (2025b). By simulating human word association tasks, we directly capture latent cultural relationships and assess the semantic spaces of cultural preferences within LLMs themselves. Extending beyond existing word association datasets, our work adapts these implicit, word-level cultural patterns to the evaluation of LLMs. Our contributions can be summarized as follows:
@@ 2 Related Work
Extensive research has investigated cultural biases in LLMs across domains such as value preference Benkler et al. (2023); Cao et al. (2023a); Jiang et al. (2024); Durmus et al. (2023), knowledge perception Palta and Rudinger (2023); Shen et al. (2024), and moral measurement Jinnai (2024); Rao et al. (2023, 2024), revealing persistent Western-centric value preferences and regionally divergent commonsense knowledge. To improve cultural perception, current approaches predominantly employ prompt engineering Wang et al. (2024); AlKhamissi et al. (2024); Zhou et al. (2025a); Masoud et al. (2025), explicitly embedding cultural context in prompts to evoke culturally sensitive responses. Alternatively, fine-tuning methods such as CultureBank Shi et al. (2024), CultureLLM Li et al. (2024b), CultureSPA Xu et al. (2024) and SimLLMCultureDist Cao et al. (2025a) have been proposed to enhance cultural alignment in different ways. However, these methods rely on survey data (particularly from word value surveys), data-driven SFT, and still require culture-specific prompts for downstream tasks.
@@
Prior to the LLM era, previous studies have also revealed cultural preferences from the perspectives of word usage and associations Garimella et al. (2016), and have attempted to use models, such as the Composite Skip-gram Model (C-SGM) Garimella et al. (2017), to predict the word association process.
On the other hand, recent work adopts word association tasks to evaluate the capabilities of LLMs across various domains. For instance, uncovering gender stereotypes Abramski et al. (2024b) and assessing color associative perception Fukushima et al. (2024). These studies primarily focus on word association mappings under specific relationships, emphasizing explicit word generation and partially relying on manual interpretation during evaluation Abramski et al. (2024a); Vintar et al. (2024).
In contrast, we not only extend word association datasets to four cultures, going beyond the two cultures (Chinese and English) which is focused on in the pre-LLM era, but also propose metrics that focus on the predicted probabilities of association words without manual intervention, enabling a more comprehensive evaluation of LLMs’ associative capabilities.
@@ 3.1 Human-centered task
The WAT for human participants is conceptualized as an open-ended activity in a free-relation format, allowing for flexible and diverse associations that more accurately reflect human cognitive processes. [^2] https://smallworldofwords.org Before conducting the WAT, demographic information, such as age, gender, education, native language, and region, is collected for each participant. Given a cue word $w$, the participant $u$ is asked to provide the first word that comes to mind and may optionally add a second and third word sequentially, resulting in up to three associated words, represented as $u_{w}^{c}=\left\{a_{1}^{u},a_{2}^{u},a_{3}^{u}\right\}$, where $c$ denotes the participant’s cultural background. Within this cultural context, the collective set of associative words for a given cue word $w$, aggregated across all participants $U_{w}^{c}=\left\{u_{w}^{c}\right\}$, is represented as $A_{w}^{c}=\left\{\left(a_{1},f_{1}\right),\cdots,\left(a_{n},f_{n}\right)\right\}$. Here, $f_{i}\geqslant f_{i+1}$, and $f_{i}$ represents the frequency with which the associative word $a_{i}$ appears in $U_{w}^{c}$.
@@ 3.2 LLM-adaptive task
Directly employing the human-centered task paradigm to conduct WAT with LLMs proves insufficient for effectively capturing variations across entire cultural groups. Inspired by Fierro et al. (2024) and Zhou et al. (2025a), we adapt this task into a word prediction framework, designed as a retrieval process for candidate associative words. Specifically, we employ a template $q$, such as “When {cue_word} is mentioned, people often associate it with the following words:”, to prompt the LLM and elicit its predictions of associated words. The input is represented as $x=q(w)$, and the output is denoted as $o=M(q(w))$, where the generalized function $M(\cdot)$ represents the inference process of the LLM.
@@ 3.3 Evaluation metric
$\begin{aligned} \text{R@K}&={\frac{\sum_{i=1}^{K}{\mathbb{I}\left(\hat{a}_{i}\in A_{w}^{c}\right)}}{N}}\\ &=\frac{\sum_{i=1}^{N}{\mathbb{I}\left(a_{i}\in\text{Top-K}\left(\hat{A}_{w}^{c}\right)\right)}}{N}\end{aligned}$ (2)
@@
A detailed comparison with other position-weighted metrics such as DCG@K is provided in Appendix B.
@@ 4 CultureSteer: Modeling and Controlling Cultural Awareness in LLMs
Inspired by the findings of LM-Steer Han et al. (2024), which suggest that, with certain assumptions, shifting styles in language models is equivalent to a linear transformation in the word embedding space, we extend this perspective to the cultural domain. Specifically, we hypothesize that different cultures define distinct semantic association spaces shaped by cultural preferences, which can be modeled through culture-specific linear transformations within the embedding space. Based on this, we introduce the CultureSteer model, which leverages both human-centered and LLM-adaptive WAT tasks to (1) equip LLMs with human-like word association perception, thereby enhancing their cognitive abilities, and (2) enable the word association capabilities of LLMs to align across different cultures. The overall architecture of the model is illustrated in Figure 2.
@@
PWR@K K=3 K=5 K=10 K=20 K=3 K=5 K=10 K=20
[…]
CultureLLM 15.32 19.89 25.10 27.76 6.01 7.48 10.34 16.15
SimLLMCultureDist 22.30 30.94 42.20 50.73 16.83 21.89 30.84 42.05
CultureMerge 20.29 29.24 40.33 49.07 - - - -
CultureSPA 19.30 27.73 40.93 49.71 - - - -
CSP 30.30 37.61 48.95 60.95 20.34 28.19 39.05 50.65
CCT 25.92 32.98 44.05 55.54 22.59 30.51 42.31 54.76
USA CultureSteer 38.14 50.30 67.58 80.13 36.34 46.06 59.78 72.05
[…]
CultureLLM 12.68 16.33 20.65 23.19 5.68 7.14 9.46 14.06
SimLLMCultureDist 19.24 26.28 35.00 41.43 14.90 19.27 26.04 34.53
CultureMerge 17.34 24.95 33.58 41.04 - - - -
CultureSPA 16.46 23.43 33.53 40.96 - - - -
CSP 24.31 30.64 39.91 49.27 16.43 22.14 30.58 39.86
CCT 21.58 27.74 35.70 44.96 18.17 24.27 33.54 42.72
UK CultureSteer 29.19 38.84 51.45 61.88 28.23 34.88 46.10 57.53
[…]
CultureLLM 13.13 16.82 20.29 22.11 5.15 5.97 8.25 13.45
SimLLMCultureDist 16.63 22.67 31.10 37.80 13.16 17.45 24.10 32.88
CultureMerge 15.99 22.64 30.64 37.83 - - - -
CultureSPA 14.42 21.20 31.35 37.78 - - - -
CSP 23.17 28.62 36.78 45.88 15.58 21.21 29.02 37.60
CCT 20.21 25.64 33.37 42.41 17.16 22.67 31.61 40.63
OC CultureSteer 27.56 35.44 47.65 58.77 25.05 32.22 41.55 52.22
[…]
CultureLLM 0.73 0.86 1.01 1.05 4.20 6.87 10.56 14.61
SimLLMCultureDist 8.42 13.33 19.09 24.72 11.78 16.32 23.82 34.98
CultureMerge 10.30 15.11 21.45 25.77 - - - -
CultureSPA 10.10 14.75 20.57 26.35 - - - -
CSP 10.52 15.83 23.34 29.81 15.56 21.59 31.77 42.81
CCT 9.71 14.08 21.02 28.00 12.32 17.71 26.57 36.85
CN CultureSteer 12.77 19.04 27.26 33.53 21.78 30.13 44.88 58.76
[…]
CultureLLM 10.47 13.48 16.76 18.53 5.26 6.87 9.65 14.57
SimLLMCultureDist 16.65 23.31 31.85 38.67 14.17 18.73 26.20 36.11
CultureMerge 15.98 22.99 31.50 38.43 - - - -
CultureSPA 15.07 21.78 31.60 38.70 - - - -
CSP 22.08 28.18 37.25 46.48 16.98 23.28 32.61 42.73
CCT 19.36 25.11 33.54 42.73 17.56 23.79 33.51 43.74
Average CultureSteer 26.92 35.91 48.49 58.58 27.85 35.82 48.08 60.14
@@ 5.1 Dataset Creation
We utilized the officially released datasets from the Small World of Words Project (SWOW) [^5] Designed as a human-centered Word Association Task (WAT).. For our study, we selected data in two languages: English and Mandarin Chinese, referred to as SWOW-EN De Deyne et al. (2019) and SWOW-ZH Li et al. (2024a), respectively. The SWOW-EN dataset comprises over 3 million responses collected from more than 90,000 participants across various English-speaking countries or regions, covering more than 12,000 cues. Moreover, the SWOW-ZH dataset contains over 2 million responses corresponding to 10,192 cues, with participants primarily from the Chinese mainland.
@@
To ensure sufficient participant representation across different cultures, we first filter the three most prominent cultural groups from SWOW-EN: the United States (USA), the United Kingdom (UK), and Oceania (OC) [^6] OC includes participant responses from Australia and New Zealand, which are combined due to their relatively participant numbers and shared historical background.. Additionally, SWOW-ZH serves as the representative for China (CN). As a result, we construct datasets representing four distinct cultural groups: USA, UK, OC, and CN.
@@
To ensure fairness in evaluating word association perception across different cultures, we utilize the Intercontinental Dictionary Series (IDS) jung Yu and Wang (2023) to align the cue words between English and Chinese, ultimately retaining 881 cue words that are common across all four cultural groups. Given the substantial differences in participant numbers across cultures, which lead to significant variation in the number of associated words ($A_{w}^{c}$), we standardize the $A_{w}^{c}$ lists by truncating them to the size of the culture with the fewest associated words. Using the IDS framework, we categorize these cue words into 22 semantic categories based on their conceptual meanings. Examples of the final preprocessed data and statistical details are provided in Appendix C.2, Table 4 and Table 5. Finally, we split the cue words in each semantic category into 70% for training and 30% for testing.
@@ 5.2 Experimental Setup
Figure 3: Fine-grained performance comparison across 22 semantic classes in the test set with PWR@20. Red denotes Global Knowledge, green denotes Perceptual Experience, and blue denotes Cultural Ideologies. Other PWR@K (K=3, 5, 10) results are shown in Appendix E.
We conduct experiments using Llama3.1-8B Grattafiori et al. (2024) and Qwen2.5-7B Team (2024) to simulate the WAT. First, we evaluate the base versions of these models on word association tasks using a base prompt without any cultural conditioning, aiming to assess their human-like associative cognition and the default cultural representation they exhibit. Subsequently, we build CultureSteer based on these two models to enhance their cross-cultural perception capabilities. For the LLM backbone, we adopt LoRA Hu et al. (2022) and train the models on a single A100 GPU.
Furthermore, to comprehensively evaluate the effectiveness of our approach, we employ both prompting-based strategies and cultural-related LLMs as baseline methods for comparison.1) Explicit Culture-Aware Prompting Strategies: We implement two prompting approaches to examine how explicit cultural conditioning affects model performance: First, Culture-Specific Prompt (CSP) Xu et al. (2025) and Cross-Culture Think (CCT). Detailed prompt templates and settings for all approaches are provided in Appendix C.1. 2) Cultural-Related LLMs: To establish comprehensive baselines, we compare against four cultural-aware language models. These include CultureSPA, Culture-merge [^7] https://huggingface.co/surbhi21/llama3-8b-cultural-merged, SimLLMCultureDist Cao et al. (2025b) and CultureLLM Li et al. (2024b) [^8] Originally trained on LLaMA2-70B, we migrate the approach to LLaMA 3.1-8B to ensure fair comparison..
5.3 Overall Results
LLMs Exhibit Cultural Biases in the WAT Task
The overall comparison results are presented in Table 1, where the “Baseline” denotes the default cultural preferences exhibited by the LLMs. Both models consistently reveal a cultural preference pattern: USA ¿ UK ¿ OC ¿ CN. [^9] To demonstrate that performance differences are due to cultural factors rather than the language used in prompts, we conducted an ablation study. Details are provided in Appendix D. This suggests that, in terms of word-level cognitive abilities, LLMs exhibit a stronger inclination toward American culture, aligning with findings from prior studies Cao et al. (2023b); Myung et al. (2024); Zhou et al. (2025a). Overall, Llama demonstrates stronger human-like word-level cognitive abilities compared to Qwen, primarily due to its significantly better performance in English-speaking cultural contexts. In contrast, the performance gap between the two models is relatively small in CN. Interestingly, when evaluating PWR@3, Qwen outperforms Llama across all cultural contexts, suggesting that Qwen excels at predicting highly associated words. However, as K increases, Llama surpasses Qwen, indicating that Llama possesses a more comprehensive associative cognition across a broader range of related words.
CultureSteer Enhance LLM Cultural Awareness
At the methodological level, CultureSteer consistently outperforms both prompt-based methods and culture-related models across all cultural settings while some models performed worse after fine-tuning. This demonstrates that the model does not rely on explicit textual configurations but also showcases the superiority of our implicit cultural steering mechanism over traditional fine-tuning approaches that may suffer from overfitting or cultural knowledge interference. Interestingly, for the Llama model, the CSP method consistently outperforms the CCT method across all cultural contexts. In contrast, for the Qwen model, the CSP method performs better only in the Chinese cultural context, while the CCT method excels in the other three, particularly in English-speaking settings. This suggests that Llama is more effective with culture-specific prompts tailored to a single context, whereas Qwen demonstrates stronger cross-cultural reasoning, especially for the English-speaking environments.